OpenStudy (anonymous):

(a + 1 - 6/a) / (a - 1 - 2/a) simplify

OpenStudy (anonymous):

lol...isnt it a^2+a-6/a^2-a-2

OpenStudy (anonymous):

\[\frac{1-\frac{6}{a}+a}{-1-\frac{2}{a}+a} \] Is this the problem expression?

OpenStudy (anonymous):

yes only with terms rearranged

OpenStudy (anonymous):

Combine the terms in the Numerator and the Denominator into one fraction respectively. \[\frac{\frac{-6+a+a^2}{a}}{\frac{-2-a+a^2}{a}} \] Invert the denominator and multiply the result by the Numerator. \[\frac{-6+a+a^2}{-2-a+a^2} \] Both the Numerator and Denominator can be factored. \[\frac{((-2+a) (3+a))}{((-2+a) (1+a))} \]\[\frac{3+a}{1+a} \] Using Mathematica for the symbol manipulation and that is why the term ordering does't look right.

OpenStudy (anonymous):

ok out casgirl

OpenStudy (anonymous):

outcastgirl